Find an equation of the line containing the given pair of points.
(-2,-6) and (-7,-9)
THe equation of the line is y= ______

Answer by KnightOwlTutor(293)

You first need to determine the slope of the line m
This is equal to m=y2-y1/x2-x1 rise/run -9-(-6)/-7-(-2)=-3/-5 3/5
y=3/5x+b
Plug in one of the points to determine b=constant
-9=3/5(-7)+b
b=-24/5
The equation is y=3/5X-24/5
To double check the answer plug in the x and y values of the second point
-6=-2(3/5)-24/5